by TEI TEI

08/29/08 – Triton is still working on warranty issues. Boiler control boards. Humidity issues. RCP problems.

02/25/08 – Triton and JCI have been working to remedy the remaining warranty issues on this project which include boiler flame outs, VAV box control, min airflows through VAV boxes, winter heating. TEI has to look at cold restrooms in the 1956 building for Kevin Maue.

12/17/07 – JCI and Triton had to go back to the site today to continue to work on DDC issues since the RTU-2 is still not running properly. Triton put the unit in had on Friday at 40%. TEI does not have an update on Monday from the troubleshooting field trip.

12/14/07 – TTC was on site today with JCI and Carrier going over warranty issues on the job including the LON talk interface between RTU-2 and DDC. Scott fixed the misnomers on the RTU-1 unit in the points list. Scott fixed the economizer program so RTU-1 would actually run in economizer mode when conditions were right (there was a programming glitch in the system) and he addressed the setback temperatures which are correct as programmed. Training will be schedule the second week of January 2008 with MRDD. There was a cold call in the lobby today from Wayne Mckee and Triton and JCI was dispatched to investigate the problems with the space. The hot water is 156F so there should be plenty of hot water available for the system.

12/04/07 – CDFP is scheduled the FM200 training directly with the MRDD folks and they will keep TTC in the loop. CDFP sent a change order in today for additional services on this project that TTC forwarded to ZZ for approval on this project for the extra sprinkler work, additional fan test.

11/28/07 – Jim Lape emailed to ask about closeout on this job. TEI stated that TJ should be 100% complete on this project. Triton is also done but we don’t have final closeout documents from them on this job. TTC is trying to get Grady to give us all closeout docs for this job.

10/17/07 – TTC visited the site and produced some more items for JCI and Triton to take care of on this job. He emailed pics and instructions to the contractors for completion. Everyone is working on completing all the work this week. Triton says the final balance reports are complete. TTC said today that he would adjust the manual high limit controllers to 210F and let the factory 190F shut the boilers down on auto reset. TTC also said he would caulk up the hole in the concrete hatch patch while he was on site. TTC will also check the installation of the electric heaters in the stairwell today while he is on site.

10/08/07 – TTC answered Jim Lape’s email about door control hardware. On drawing E400 #12 & #13 state that TJ will wire up the devices furnished & installed by the GC.

10/01/07 – TTC was on site last week performing commissioning. The control system was not ready to be back checked by TEI. JCI had not performed the functional test and TTC did not have the CX sheets. Nevertheless the system is running and with another week of actual work, it will be fine. TTC requested trend reports for verification of working systems.

09/19/07 – TTC approved JMW change order for roof drain and sent email to Jim, Zach and Kevin to that effect. TTC reviewed Triton invoice and asked for change order modifications before actual billing to the County. Change Orders were not reflected on actual payapp but they were on the schedule of values.

09/17/07 – Discussed Change Orders # 09,10,11 with Tim and Kevin Maue. I approved them as submitted. The steel issue is still are large issue that doesn’t appear close to resolution. TTC responded by email about controlling the pressure of the system by setting the static residual to 5 psig and letting the building set the pressure from there. If we have to change the relief valves to 65 psig from there, we can. The static of the building on the basement pumps should be 26.64 psig (50 ft bldg ht / .4328 ft h20/ psi + 5 psig to keep air out).

09/12/07 – Triton is on site checking the existing VAV boxes for repairs. They found one bad damper and believe 6 of the 7 boxes that are not operating properly. We are concerned that we have paid for some of this work already at the beginning of the job, but we will pay for the repairs to the dampers now and fix the electric heaters.

09/04/07 – JCI has given training but the controls are not complete and they still need some work. Boiler are not operational at this time and Denny Lammers is still need to program the internals of the boilers so JCI can communicate and give them their firings rates and monitor alarms. TTC dropped off TJ Williams hard copy as-builts drawings for the project that need to be included into the O&M’s that Kevin already has in his office.

08/23/07 – TEI printed out 2 more originals for Change Order #1, signed them, and sent them to T.J. Williams with instruction to sign them and forward them to HCFD.

08/17/07 – TEI received as-built drawings from TJ. Williams today. TEI will update the electronic drawings with the information and forward them back to CRA early next week. TTC delivered the O&M’s for EC, PC and FPC to Jim Lape at his office today.

08/16/07 – TTC and Cintas performed and passed the final server room fire alarm test on Wednesday. TTC has the paperwork in the folder and will include in the O&M. Triton said that the handrail was ordered and will be put back by Pat Murphy.

08/06/07 – TJ will do the security camera and emergency lighting change order work at T&M not to exceed but they will not come down in price. TTC sent this info to Jim Lape and approved on his end. TEI needs direction from the owner. TTC contacted MRDD about installing gaskets behind wall plates in server room.

08/02/07 – TTC talked to Kevin today about TJ outstanding issues at the MRDD. Security Camera change order and Emergency Lighting change order. JMW has O&M’s into TEI and they all closeout paperwork into Kevin Maue. TEI received TJ Williams closeout paperwork on this project as well. NV had TTC give to it to NR for approval.

07/23/07 – TTC spoke to Tim Myers today about project closeout, change orders, punchlist and controls. TTC discussed the allowance usage with Tim which totaled $13459 to date not including the refrigerant changeout which is a different cost. TTC emailed CRA about distributing the correct paperwork for change order pricing from Triton on this job and listed 6 items that need to be priced for this project.

1. Extra Diffuser for 1956 building. 17 in Qty. Linear Slots and 2×2’s to meet job requirements.
2. Refrigerant Check, Removal, Replacement on RTU-1. This work wasn’t included in the allowance cost but we authorized Triton to do the work on a separate ticket.
3. Overtime for IT closets and occupation over one weekend (i believe KM authorized triton to do this work)
4. Replacement cost for handrail? Triton removed it but someone else threw it in the garbage. Hobert is supposedly looking for it today but Triton will have to replace it if it can’t be found.
5. Demolition of all housekeeping pads in boiler room. This was Empire work but Triton had to do it all because Empire refused to do it. If Triton is going to pay the steel cost as a change order to Empire which I now believe they will after speaking with Tim Myers this morning, Empire either has to reimburse them this cost or deduct from the steel change order cost. Tim brought up a good point. A $2029 redesign fee is steep. That is 20 hours of a ThermalTech Principal time or 40 hours of a contractors time. It did not take that long. Redrawing that steel took about 6 to 8 hours. That fee should from the sub-contractor $500-$700. Please try to get them to give us a “real” cost and the housekeeping pad back charge could go away. As a side note, the new designed decreased the amount of steel and grating required, why are they paying an upcharge of $2522.97?
6. Replacement of 2 electric heaters. Provide 2 recessed ceiling heaters in the 1956 stairwell in lieu of the scheduled heaters. Return the surface mounted fin heater and the wall heater.
07/17/07 – TTC has scheduled a backcheck of the plumbing punchlist on Thursday. He will meet Jason on site for the walk through. Triton has committed to Approved Products being on site to finish the balance portion of the work starting on 07/23. TTC sent 2 sketches to CRA for plumbing ASI’s.

07/03/07 – The job is mostly complete aside from outstanding issues. The MRDD people are in and using the building. Items that we are still working on include: balancing HVAC & water, relocating a fp line out of the elevator shaft, final test of FM200, hot water boiler loop, controls, roof drain near unit, and punchlist by all contractors.

06/26/07 – All TEI punchlist were generated last Friday. TEI gave Jim Lape a complete set of extra materials required on this project to included in substantial completion for the project. TTC was onsite with JMW plumbing early Monday morning to discuss punchlist items which they are addressing. TTC spoke with Kevin about roof drain issue on 1956 building and we agreed that changing the scupper might be the best option. We can drain it down to the canopy roof.

06/15/07 – TE will be generating a punch list next week.

06/07/07 – RTU is still surging. we are trying to find out the status of the jci & carrier contractors. JMW plumbing will be on site on Friday to demo the 4″ plumbing vent from the elevator room. we believe fire protection is complete and ready for testing.

06/04/07 – Inspections are scheduled for Tuesday 06/05. MRDD personnel are moving equipment into the building today. ttc did not approve payapp #6 for triton and asked them to back down the % on the ground floor of the 1956 bldg to 90% and controls down a bit lower.

05/30/07 – TEI was on site on Tuesday for our weekly walk through. TEI contacted the vendors with questions, TCO milestones and coordination issues at the MRDD to help get it ready for the MRDD move on June 4. Items include ceiling layout, diffuser locations, ductwork insulation, EF flashing, fire protection work, elevator details, sump work, controls, etc.

05/21/07 – Triton reported that all fire dampers were completed over the weekend and that mason would be finishing up wall penetrations today and tomorrow. Kevin Maue seemed unimpressed with the quality of the work being done over the weekend and hopes they pass inspection. Tim, Terry and Kevin will walk the site on Tuesday AM to check progress and get commitments for completion.

05/17/07 – Triton will not be working this weekend but will have another crew on the MRDD job starting on Monday and they will have masonry sub helping with the patching of the duct shaft. The new RTU will not be ready to be ran this weekend.

05/14/07 – A lot has happened since my last entry. We have attended weekly const mtgs. We have attended several additional field trips for construction issues. The 1992 building has reached occupancy. The 2nd floor move was today. The 1992 RTU-1 unit is up and running but not under full control. The FM200 system is not fully tested yet even though the room is built. The 1956 building has ceiling grid in the entire building and now diffusers and fire protection piping is being installed.

04/17/07 – TTC and TA walked the project with all the contractors in lieu of a const mtg. The RTU is on the roof. The Server room is nearly complete. Cintas is working to complete the testing procedure and set it up for us. Shaft wall is offset form fire dampers. TE approved shifting the wall over to catch the sleeve of the fire dampers. The wall holes in the return shaft are open and Triton asked about filling them in (even the ones that weren’t done by them). TE directed Triton that they could sleeve the chimney with the boiler vents. The elevator room exhaust fan will reuse one of the openings that goes up to the chimney. A gooseneck may have to be added to the top of the chimney to keep the exhaust 10ft away from our unit intake.

04/13/07 – Project progressing as expected.

04/06/07 – Project progressing as expected.

03/30/07 – Project progressing as expected. Triton reported that the Liebert unit can’t be started today because of a breaker size issue (req 40A we have a 20A spec’d). TA resolved with TJ and Triton will have startup technician for unit after power is provided. TE received submittals today from JCI and GD has been looking them over all day. KM & TTC emailed Triton with acceptance on the retro-commissioning items on RTU-1.

Replace the Supply Fan Motor $3566.00 – APPROVED

Replace two bad compressor contactors $408.00 – APPROVED

Repair static pressure control for exhaust $2720.00 – APPROVED

Replace one bad condenser fan motor $1149.00 – APPROVED OR REPLACE WITH ALL 6 DEPENDING ON PRICE TBD

Perform leak check of Ref system, repair and replace refrigerant as needed $2339.00 – HOLD

Cleaning of coils and inside if unit including combing of coils as needed $1633.00 – APPROVED

Repair of Inlet Guide Vane and controls $3931.00 or replace same with VFD control $5108.00. – APPROVED BASED ON AVAILABILITY OF PARTS

03/23/07 – Lobby lights are on order and all issues with this item have been resolved. 480V electrical service has been changed over by Duke Energy, has been inspected and is up and running. 240V service to be removed the last week of March 07, tentatively.

03/20/07 – 02/19/07 – Last week the structural steel modifications began for the Carrier unit changes. Triton provided a dwg to SSA and CRA showing the new steel dims. SSA is sending back a proposal. The return duct shaft is open and must be sealed to provide proper airflow back into the unit. CRA is checking on some mold issues in that shaft from a roof leak. The FPC dwgs have been submitted to the city. The new FM is located and routed through an existing coredrills in the connector. TA release approval on change orders for TJ Williams for emergency lighting battery packs and dishwasher power in the 1992 building. Triton is still behind on the VAV boxes for the 1992 building and will have to work with Empire for ceiling pad out.

02/09/07 – TTC walked the site last Tuesday to check on progress. MEP is moving forward okay. Duct chase holes have not been cut yet. Triton is installing ductwork on the 1992 building

01/17/07 – Triton has coordinated with Empire on existing work. Kevin is worried that we don’t have enough written documentation that we told Triton to get this work coordinated and scheduled with the owner

01/15/07 – Somehow this job has not been tracked. Here’s the catch up data. The MC is Triton, the EC is TJ Williams, the PC JMW plumbing. We have had several job meetings. Triton has been poorly represented at these meetings. Last week after missing a rescheduled job meeting, Triton came on board with some initiative in catching up to the rest of the contractors. Triton has added more people to the job and is beginning new HVAC in the 92 building. Triton has demolished two floors of work in the 56 building. Triton has 2 men finishing the demo of the multi-zone unit and boilers in the basement of the 56 building. Triton has submitted an RFI questions who is cutting the holes in the shaft floor for his ductwork. TTC responded the GC and this was confirmed by CRA. The RFI was replied to but Empire responded they did not agree. The work is shown on the Demo and Arch dwgs for the GC and the summary of work says superstructure work is by the GC. The CWW has denied the underground permit and tap because the CGE pit is in the Right of Way of the street. TEI has negotiated with them and Cin-Day Fire Protection to provide the new line inside the building through the connector and throughout the building. This should save money and remove trenching and such in the yard. Kevin Maue has agreed that he likes this approach.

09/14/06 – Weekly design meetings continue. We reviewed cost estimates with the client and architect. Final estimates and drawings are due on September 25th for review and bidding is scheduled shortly thereafter.